OpenStudy (anonymous):

someone please help me!! Find the GCF of the monomials: 18x2, 30x2y, and 54y 18 6 18xy 6xy

OpenStudy (anonymous):

well it cant have an x or y because not all of them have x and y so now you just have to figure out the greatest number that goes into all 3 of the numbers

OpenStudy (anonymous):

i mean i just dont understand itwhat it wants me to do?

OpenStudy (anonymous):

find the biggest number that goes into all 3 of the numbers

OpenStudy (anonymous):

80,30, ande 54?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2

OpenStudy (anonymous):

sorry it is 18* 30 and 54 and look at your answer choices. then try it

OpenStudy (anonymous):

how do i find the biuggest number that goe sinto all of them?

OpenStudy (anonymous):

take each number and divide it by 6, then if all of that works, try 18

OpenStudy (anonymous):

\[18=2\times 3^2, 30=2\times 3\times 5,54=2\times 3^3\] greatest common factor is \[2\times 3=6\]

OpenStudy (anonymous):

or you could do it that way...

OpenStudy (anonymous):

i got 18= 2,3,6,9 30= 2,3,5,6,10 54= 2,27 so is it 2?

OpenStudy (anonymous):

@satellite73 is right. for 54 you didnt get all of the factors

OpenStudy (anonymous):

i made a mistake the gfc is 6 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the answer is b?

OpenStudy (anonymous):

yesh

OpenStudy (anonymous):

thanks!